OpenStudy (anonymous):

3x + 2y = -4 x + 2y = -8 Solve the systems using substitution

OpenStudy (anonymous):

3x + 2y = -4 =>3(-2y-8) + 2y = -4 =>-6y-24+2y=-4 =>4y=-20 =>y=-5 =>x=2

OpenStudy (anonymous):

from the second equation we get x= -8 -2y substituting in first 3(-8-2y) + 2y = -4 -24 -6y + 2y = -4 4y = -20 y = -5